OpenStudy (anonymous):

Find the derivative and simplify: f(x)= [x(x^2-1)]/[x+3] I'm not sure if I should use the Product Rule, the Chain Rule, the Quotient Rule, or a combination of two!! I would love some help and 'thank you' in advance!

OpenStudy (phi):

looks like you use all of the above

OpenStudy (anonymous):

I looked at the work I did for this and I honestly don't think I did it correctly. I believe I used the product rule and then the quotient rule? I got something like this: f'(x)= (2x^3+9x^2-3)/(x^2+9x+9)

OpenStudy (phi):

I would not expand the factors... you could start with the quotient rule \[ d \frac{u}{v}= \frac{v\ du - u\ dv }{v^2} \] where u is x(x^2-1) and v is (x+3) to find du you use the product rule d( a b) = a db + b da where x will be a and (x^2+1) is b to find db you find d(x^2 + 1) and use the power rule.

OpenStudy (anonymous):

so you think I should do the quotient rule and then the product rule? I think I did it in the opposite order. I'll try working this out, then.

OpenStudy (anonymous):

But if my earlier answer was correct, you meant to not expand and keep it like this: (2x^3 +9x^2 -3)/(x+3)^2 ?

OpenStudy (phi):

yes, I would not expand that out unless necessary (for example, multiple choice had the answers in expanded form)

OpenStudy (phi):

also, for this problem, it might be easier to expand x(x^2 -1) to x^3 - x so that you don't have to use the product rule. (or do it both ways, as a check. You should get the same answer)

OpenStudy (anonymous):

ok, I'll try that now to check my work!

OpenStudy (anonymous):

ok, so I was correct! yay! I was just a little unsure of myself. Thank you for your help! I'll award Best Response if I can, but I'm not sure if I have any medals...

OpenStudy (phi):

btw, your answer up above **** I got something like this: f'(x)= (2x^3+9x^2-3)/(x^2+9x+9) ***** looks ok, except the bottom should be (x+3)^2 = x^2 +6x+9 It looks like you should expand the terms up top, because it simplifies down quite a bit.

OpenStudy (anonymous):

oh ok so if I have (2x^3 +9x^2 -3)/(x^2+6x+9) I didn't see any way to simplify it...? no common factors as far as I can tell

OpenStudy (phi):

that looks good.

OpenStudy (anonymous):

ok then! thank you for your time and help! i'll close this question.

OpenStudy (phi):

I was thinking that after we take the derivative, we could leave the top "as is", but it makes sense to do what you did, and rewrite it as 2x^3 +9x^2 -3

OpenStudy (anonymous):

ok that's what I had thought. my calc teacher likes things simplified, but not too simplified. :3

OpenStudy (phi):

though the bottom I would leave as (x+3)^2 it looks nicer that way

OpenStudy (anonymous):

alright, well I have what you just said as the answer and I have it expanded underneath, just in case!